OpenStudy (anonymous):

Thanks for helping! Help with symmetries! Identify, graph, and state the symmetries of each polar equation. Write the scale that you are using for the polar axis. r=9cos(5 theta)

OpenStudy (anonymous):

@Hero @jim_thompson5910 @phi @zepdrix @SithsAndGiggles @wio @sleepyjess @Data_LG2

OpenStudy (anonymous):

I can't figure out the symmetries though

OpenStudy (anonymous):

 You can check if its symmetry by putting in \(-\theta\) in \(\theta\) \(r=9cos(5\theta)\) Replacing it, \(r=9cos(5(-\theta))\) \(r=9cos(-5\theta)\) Can be written as \(9cos(5\theta)=r\) Which means this equation is symmetrical